OpenStudy (anonymous):

the perimeter of a rectangle is 256 cm. if the sides are in the ratio 2:5. find the rectangle?.. with medal

OpenStudy (ranga):

Let L be the length and W be the width of the rectangle. Perimeter = 2(L+W) = 256 divide both sides by 2 L + W = 128 ----- (1) Sides are in the ratio 2:5 That means W/L = 2/5 cross multiply 5W = 2L divide both sides by 5 W = 2L/5 ---------- (2) Put W from equation (2) into (1) and solve for L. Then put the L value in (2) and solve for W.

OpenStudy (anonymous):

@ranga what??

OpenStudy (anonymous):

can you explain it clearly

OpenStudy (ranga):

Do you want to know how to solve or do you want to know how I got the 2 equations?

OpenStudy (anonymous):

i want to know how to solve

OpenStudy (ranga):

L + W = 128 ----- (1) W = 2L/5 ------ (2) Put W = 2L/5 in the first equation L + 2L/5 = 128 multiply each term by 5 to get rid of the fraction: 5L + 2L = 640 7L = 640 L = 640/7 = 91.43 cm Put this in (2) W = 2L/5 = 2*91.43/5 = 36.57 cm. L = 91.43 cm; W = 36.57 cm.